OpenStudy (anonymous):

Which of the following is a true statement? All square roots are irrational. All decimals are irrational. All non-perfect squares have irrational square roots. All irrational numbers can be written as ratios of two integers.

jimthompson5910 (jim_thompson5910):

which one do you think it is

OpenStudy (anonymous):

this one All square roots are irrational.

jimthompson5910 (jim_thompson5910):

what about sqrt(25) ?

jimthompson5910 (jim_thompson5910):

is that irrational?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wow you have alot of people veiwing

jimthompson5910 (jim_thompson5910):

you are correct, sqrt(25) = 5 = 5/1, which shows it's rational

jimthompson5910 (jim_thompson5910):

since it can be written as a fraction of whole numbers

OpenStudy (anonymous):

so b then

OpenStudy (anonymous):

so is this the correct answer All decimals are irrational.

OpenStudy (anonymous):

im still her courtney

OpenStudy (anonymous):

yess because 0.2145214 cant be written as a faction

jimthompson5910 (jim_thompson5910):

no, 2.5 is rational since 2.5 = 25/10 = 5/2

OpenStudy (anonymous):

no you just dont get rid of the demical

jimthompson5910 (jim_thompson5910):

but the fact that 2.5 = 5/2 means that 2.5 is rational

OpenStudy (anonymous):

listen badgurl it is B

OpenStudy (anonymous):

ok jim wat do u think it is

jimthompson5910 (jim_thompson5910):

All non-perfect squares have irrational square roots.

jimthompson5910 (jim_thompson5910):

since sqrt(25) = 5, but 25 is a perfect square however, sqrt(26) is irrational since you can't write it as a fraction of two integers notice how 26 is a non-perfect square

OpenStudy (anonymous):

oh ok yeah i remember my teach talking about that

OpenStudy (anonymous):

you know what just go with the one you think it is i want be mad

OpenStudy (anonymous):

ok courtney and jim i have another one

OpenStudy (anonymous):

so you want me to leave

OpenStudy (anonymous):

Approximate \[\sqrt{110}\]to the nearest tenth without using your calculator.

OpenStudy (anonymous):

do you want me to leave or what

OpenStudy (anonymous):

A. 10.0 B. 10.1 C. 10.4 D. 10.5

OpenStudy (anonymous):

and no i dnt want you to leave help me

OpenStudy (anonymous):

ok i say 10.4 but that is with i with the caculator

jimthompson5910 (jim_thompson5910):

10*10 = 100 11*11 = 121 So the number is between 10 and 11

jimthompson5910 (jim_thompson5910):

So let's say it was 10.5 (that's the midpoint between 10 and 11)

jimthompson5910 (jim_thompson5910):

so 110/10.5 = 10.47619

OpenStudy (anonymous):

ok i have a more ? courtney u still there

jimthompson5910 (jim_thompson5910):

notice how close we are already

OpenStudy (anonymous):

yea but 10.4 is the correct answer with i caculator

OpenStudy (anonymous):

ok next ?

OpenStudy (anonymous):

Which of the following expressions would appear furthest to the right on a number line? \[A. \sqrt{3}+5 B.7 C. \sqrt{36} D.\frac{ 15 }{ 2 }\]

OpenStudy (anonymous):

ok badgurl have to go but i will be back on in a few

OpenStudy (anonymous):

i promise

OpenStudy (anonymous):

jim help me

jimthompson5910 (jim_thompson5910):

Are you allowed to use a calculator here?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

sqrt(3) = 1.73205

jimthompson5910 (jim_thompson5910):

so sqrt(3) + 5 = 1.73205 + 5 = 6.73205

jimthompson5910 (jim_thompson5910):

do you see how I'm getting this?

OpenStudy (anonymous):

yeah a lil

jimthompson5910 (jim_thompson5910):

what is sqrt(36)

OpenStudy (anonymous):

i have no clue

jimthompson5910 (jim_thompson5910):

you agree that 6 times 6 = 36 right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so sqrt(36) = 6

jimthompson5910 (jim_thompson5910):

finally, what is 15/2

OpenStudy (anonymous):

it an improper fraction

jimthompson5910 (jim_thompson5910):

use calculator to find out what 15 divided by 2 is

OpenStudy (anonymous):

7.5

jimthompson5910 (jim_thompson5910):

so the original list of numbers becomes 6.73205, 7, 6, 7.5

jimthompson5910 (jim_thompson5910):

which is the largest one

OpenStudy (anonymous):

7.5

jimthompson5910 (jim_thompson5910):

that corresponds to 15/2

jimthompson5910 (jim_thompson5910):

so that is the largest number, so it's the furthest to the right on the number line

OpenStudy (anonymous):

convert 325 f to appropriate metric units

jimthompson5910 (jim_thompson5910):

to meters?

OpenStudy (anonymous):

i got it thanks though

jimthompson5910 (jim_thompson5910):

ok np